ANALYSES OF VIBRATIONS OF Ar_{n} AND Ar_{n}CO_{2} CLUSTERS BY MOLECULAR DYNAMICS CALCULATION

\underline{Y. OZAKI}, Department of Chemistry, Faculty of Science, Josai University, Sakado, Saitama 350-02, Japan; M. ICHIHASHI, T. KONDOW, Department of Chemistry, School of Science, The University of Tokyo, Bunkyo-ku, Tokyo 113, Japan; AND K. KUCHITSU, Department of Chemistry, Faculty of Science, Josai University, Sakado, Saitama 350-02, Japan.

Previously, we analyzed intracluster vibrational motions of Ar_{n}, n=13–55, clusters by use of molecular dynamics method ^{a}. It was shown that a number of collective vibrations were excited. The vibrations were assigned to breathing, quadruple spheroidal and quadruple torsional vibrations corresponding to the modes of a dense sphere with the radius of Ar_{n} and the phonon velocities in solid Ar. Such vibrations were observed for Ar clusters by the energy loss of colliding He atoms ^{b}. The frequencies of these vibrations were found to decrease with the increasing cluster size, suggesting that these vibrational motions are new physical properties of clusters obeying a scaling law. In this work, the motions of Ar atoms and CO_{2} molecule in Ar_{n}CO_{2} cluster produced by addition of CO_{2} to the surface of Ar_{n} cluster, are simulated by use of the classical molecular dynamics method. The potential is calculated by overlapping two-body Lennard-Jones potentials, where an anisotropic potential is used for the Ar-CO_{2} interaction. By extracting the breathing and quadruple spheroidal vibrations, the frequencies of these modes are calculated to be: e.g., 33 and 20 cm^{−1}, respectively, for Ar_{13}CO_{2}. These values are almost identical to those for Ar_{13}. This indicates that the addition of CO_{2} causes little effect on the vibration of Ar_{13} when the CO_{2} molecule is on the surface of Ar_{13}. Cluster-temperature and cluster-size dependence of the frequencies will be presented.

aY. Ozaki, M. Ichihashi and T. Kondow, Chem. Phys. Lett. 182, 57 (1991); ibid. 188, 555 (1992).

bU. Buck and R. Krohne, Phys. Rev. Lett. 73, 947 (1994).